Below is an ANOVA table. Use it to answer the following 5 questions. Analysis of Variance Source df Sum of Squares Mean SquarQuestion 15 1 pts What is MS between? Question 16 1 pts What is MS within? Question 17 1 pts What is Fobt?
Answer #1

Given

DF between = 2

Df total = 29

Now calculate Df within

1) Df within = Df total - Df between

= 29 - 2

Df within = 27

SS within = 468

SS TOTAL = 818

2) NOW Calculate SS BETWEEN

SS Between = SS TOTAL - SS WITHIN

= 818 - 468

SS BETWEEN = 350

3) MS BETWEEN = SS BETWEEN / DF BETWEEN

= 350/2

MS BETWEEN = 175

4) MS WITHIN = SS WITHIN / DF WITHIN

= 468 / 27

MS WITHIN = 17.33

5) F = MS BETWEEN / MS WITHIN

= 175/17.33

F = 10.098